• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name P.Vernin

    Log entry time 10:18:34 on October 24,1998

    Entry number 3434

    keyword=ARC energy measurement

    Arc tunned in dispersive mode, chicane off, pulsed beam, empty target:
    made a scan, could not see the beam ( bad timing?), ask for CW.
    9:40 MCC calls to say ready with CW, we ask for 5uA.
    CW: our scan trip Compton chicane BLM: ask MCC to mask it
    gains=4/16/64/64, file=oct2498-ARC: #2 is saturated
    gains=4/8/64/64, file=oct2498-1-ARC
    HV changed from -1299V to -1300V on PMT2 and gains=4/8/32/32:
    file=oct2498-3-ARC
    same as previous: file=oct2498-4-ARC
    Dipole shed measurement done, file=integral-248.data
    Second dipole shed measurement: file=integral-249.data

    The quality of the data seems good, need analysis now.